# G2SF-MIAD: Geometry-Guided Score Fusion for Multimodal Industrial Anomaly Detection

### v1

@@ IV-A Experimental Details
TABLE III: Analysis of various score metrics for anomaly detection.
@@
TABLE IV: Ablation results of individual loss components within G2SF.
IV-B Comparison on Multimodal Benchmarks
We illustrate the results of I-AUROC and AUPRO@30% in Table I, while those of AUPRO@1% and P-AUROC are presented in Table II. Specifically, for anomaly detection, the proposed G2SF achieves the highest I-AUROC value of 97.1%. Although LSFA [6] matches G2SF in terms of I-AUROC, its AUPRO@30% result for anomaly segmentation is 96.8%, which is lower than G2SF’s 97.9%. Regarding anomaly segmentation, our G2SF outperforms all competitors in three pixel metrics with a +1.2% AUPRO@1% advantage over Shape-guided [19]. Notably, for AUPRO@30% value, although Shape-guided is close to G2SF, its I-AUROC of 94.7% is significantly lower than G2SF’s 97.1%. Overall, G2SF achieves the best performance in both anomaly detection and anomaly segmentation tasks. This performance gain stems from our purpose of learning a discriminative local distance metric. As demonstrated by the qualitative results visualized in Fig. 6, G2SF effectively suppresses overestimated scores in normal regions compared to M3DM [7] and CFM [18], while producing sharper anomaly boundaries. To provide a comprehensive visualization, Fig. 7 illustrates pixel-wise anomaly score distributions in four categories, demonstrating that anomaly scores of normal regions in G2SF shrinkage significantly within a narrow range. This property effectively reduces false positives, leading to more substantial gains in AUPRO@1% (Table II). More qualitative results are provided in the supplementary material.
IV-C Analysis
Effectiveness of Final Score ${s}_{i}$: Table III compares key score metrics for anomaly detection: unimodal scores $s^{m}_{i}$ ($m\in\{P,R\}$) (Euclidean distances, see Section III-B), scaling factors $w^{m}_{i,0}$ from LSPN in Eq. (2), and our fused score ${s}_{i}$ in Eq. (11). As evidenced in Table III, unimodal anomaly scores $s^{m}_{i}$ demonstrate suboptimal performance. This limitation stems from the inherent information incompleteness of individual modalities. Our fused score ${s}_{i}$ overcomes these constraints through cross-modal complementarity, achieving significant improvements over unimodal baselines. These quantitative gains are also supported by Figs. 6 and 8.
While $w^{m}_{i,0}$ characterizes the local directions of normal features within local spaces (Section III-D), its exclusion of the valuable distances, i.e., $s^{m}_{i}$, leads to critical limitations. Specifically, proximity to memory prototypes ($s^{m}_{i}\rightarrow 0$) inherently indicates normality regardless of directional deviations, which is a prior ignored by $w^{m}_{i,0}$. As evidenced in Fig. 8, the score maps obtained by $w^{m}_{i,0}$ exhibit noticeable false positives. Using ${s}_{i}$ in our G2SF mitigates this issue, which benefits from sufficiently leveraging complementary directional and distance information. The comprehensive results of all categories are provided in the supplementary material.
@@ IV-C Analysis
TABLE V: Investigation on different score aggregation strategies.
Ablation of Individual Loss Components: Table IV systematically evaluates the contribution of each loss component through stepwise removal. Using the separation loss $\mathcal{L}_{sep}$ as the baseline for the metric learning, the addition of the consistency loss $\mathcal{L}_{cns}$ drives an 8.2% I-AUROC improvement, surpassing M3DM [7] across all metrics and achieving parity with CFM [18]. The margin loss $\mathcal{L}_{mar}$ specifically improves pixel-wise localization (AUPRO@30% +0.3%), while the scaling loss $\mathcal{L}_{sc}$ and cross-modal alignment loss $\mathcal{L}_{cma}$ further enhances image-level detection (I-AUROC +1.1/1.0%). These synergistic effects collectively establish G2SF’s state-of-the-art performance, validating the necessity of all proposed loss components.
Investigation of Score Aggregation Strategy: Section III-F employs a $\min$ operator in Eq. (11) to aggregate metrics $\{l_{i,j}\}$ from adjacent local spaces. We systematically evaluate three alternatives in Table V: (1) single-space baseline $l_{i,0}$. (2) $\max$ operator. (3) $\mathrm{mean}$ operator. The $\min$-based strategy achieves superior overall performance (+0.8%/0.7% I-AUROC/AUPRO@1% vs $l_{i,0}$), validating the necessity of multi-space consensus. While $\max$ and $\mathrm{mean}$ operators incorporate adjacent local spaces, they indiscriminately weight less discriminative $l_{i,j}$ from distant regions, compromising detection precision. The $\min$ operator in Eq. (11) prioritizes the most representative local metric through a proper geometric prior as discussed in Section III-F. The results of all categries of MVTec-3D AD are summarized in the supplementary material.
V Conclusion
In this paper, we propose G2SF for multimodal anomaly detection, which learns a unified anisotropic local distance metric to replace the indiscriminative isotropic Euclidean metrics. This metric is formulated by the direction-aware scaling factors predicted by LSPN and original Euclidean distances in individual modalities. We propose new loss functions to enable robust metric learning, serving for discrimination or geometric property maintenance. A geometric prior-induced anomaly scoring strategy is proposed to enhance anomaly detection and segmentation. Extensive experimental results show our superior performance with low positive rate and better recall compared to existing state-of-the-art methods.
@@ B Implementation Details
B-A Feature Extraction
RGB features are extracted using DINO ViT-B/8 [20] pretrained on ImageNet [31], generating $28\times 28\times 768$ feature maps upsampled to $56\times 56\times 768$. Point clouds are processed via Point-MAE [21] pretrained on ShapeNet [32], grouping 3D points into 1024 clusters (feature dimension 1152) and interpolated to $56\times 56\times 1152$ for spatial alignment with image feature maps. We use PatchCore [5] to construct modality-specific memory banks with 10% coreset sampling. This procedure remains the same as M3DM [7] without any change.
B-B Data Processing
To accelerate training and focus on discriminative foreground features, we pre-collect features from 800 synthetic anomalous samples per class (Section III-G), forming a feature dataset equally split for training and validation. The features originate from foreground regions identified via RANSAC plane fitting on 3D point clouds, where areas exceeding 0.005 distance from the fitted plane are retained following [7, 18]. During evaluation on MVTec 3D-AD, background regions remain included. As background features are excluded during training, we set their scaling factors to unity ($w^{m}_{i,0}=1$) in Eq. (3), serving as a linear combination of the original unimodal results.
B-C Parameter Setting
For the consistency loss $\mathcal{L}_{cns}$, we set $k=5$ and $\eta_{0}=1.2$. The tuning parameters associated with each loss are set as $\alpha=10,\beta=60,\gamma=8,\mu=20$. The LSPN is trained using an Adam optimizer [30] with learning rate $1.5e^{-4}$ and weight decay $1.5e^{-4}$, combined with $l_{1}$ regularization of weight $1e^{-4}$ and drop out with ratio 0.5, over 80 epochs with a batch size of 8,192. Global scaling factors $\sigma^{m}$ are separately optimized at a higher learning rate of $5e^{-3}$ to accelerate convergence. We initialize LSPN by the default strategy in Pytorch, that is, Kaiming Initialization [33]. We observe $w^{m}_{i,j}\approx 1$ under this strategy, meeting the requirement in Section III-D. For modality $m$, we normalize all Euclidean distances $\{s^{m}_{i,j}\}$ by their mean over the training dataset. Then, we initialize $\sigma^{m}=0.5$ with equal importance from each modality. Finally, we implement our G2SF by Pytorch and run all experiments on Linux with Nvidia RTX 4090 GPUs.
@@ C Learning Curve
Figure 3: Performance over different training epochs.
@@
We train multiple models that stop at {2,4,6,8,10,20,…,80} epochs respectively, and report their performances (I-AUROC, P-AUROC, AUPRO@30%, and AUPRO@1%) on MVTec-3D AD in Fig. 3. After only 4 epochs, G2SF outperforms M3DM [7] (94.5% I-AUROC, 99.2%P-AUROC, 96.4% AUPRO@30%, and 39.4%AUPRO@1%). This benefits from our design, as our model evolves from an established Euclidean metric to the target anisotropic metric, see Section III-D.
D Additional Quantitative Results
@@
We summarize the results of different score metrics for all categories in MVTec-3D AD in Table I. The results demonstrate that the final score $s_{i}$ outperforms unimodal scores $\{s^{m}_{i,0}\}$ and scaling factors $\{w^{m}_{i,0}\}$ almost all categories. Additionally, comprehensive quantitative results of various score aggregation operators are provided in II, which validate the superior performance of $\min$ operator over the majority of classes, owing to its geometric interpretation, as discussed in Section III-F.
@@
TABLE I: Results of various score metrics on MVTec-3D AD. Best results in bold, runner-ups underlined.
@@
TABLE II: Results of various score aggregation strategies on MVTec-3D AD. Best results in bold, runner-ups underlined.
@@
E Additional Qualitative Results
Fig. 4 provides additional qualitative results for all classes in MVTec-3D AD. Consistent with Fig. 6, our G2SF suppresses anomaly scores of normal regions and provides clearer anomaly boundaries, compared to M3DM [7] and CFM [18].
Figure 4: Additional qualitative results on MVTec-3D AD.

### v2

@@ 4.1 Experimental Details
Table 3: Analysis of various score metrics on MVTec-3D AD.
4.2 Comparison on Multimodal Benchmarks
We illustrate the results of I-AUROC and AUPRO@30% on MVTec 3D-AD in Table 4, while those of AUPRO@1% and P-AUROC are presented in Table 2. Specifically, for anomaly detection, the proposed G2SF achieves the highest I-AUROC value of 97.1%. Although LSFA [30] matches G2SF in terms of I-AUROC, its AUPRO@30% result for anomaly segmentation is 96.8%, which is lower than G2SF’s 97.9%. Regarding anomaly segmentation, our G2SF outperforms all competitors in three pixel metrics with a +1.2% AUPRO@1% advantage over Shape-guided [10]. Notably, for AUPRO@30% value, although Shape-guided is close to G2SF, its I-AUROC of 94.7% is significantly lower than G2SF’s 97.1%. Overall, G2SF achieves the best performance in both anomaly detection and anomaly segmentation tasks. This performance gain stems from our purpose of learning a discriminative local distance metric. As demonstrated by the qualitative results visualized in Fig. 5, G2SF effectively suppresses overestimated scores in normal regions compared to M3DM [32] and CFM [11], while producing sharper anomaly boundaries. To provide a comprehensive visualization, Fig. 6 illustrates pixel-wise anomaly score distributions in four categories, demonstrating that anomaly scores of normal regions in G2SF shrinkage significantly within a narrow range. This property effectively reduces false positives, leading to more substantial gains in AUPRO@1% (Table 2). In addition, the results on Eyecandies in Table 2 showcase the best performance of our G2SF over all metrics, particularly with +2.2% AUPRO@1% over CFM [19]. More details are provided in the supplementary material.
4.3 Analysis
@@
Table 4: Ablation results of individual loss components within G2SF on MVTec-3D AD dataset.
Effectiveness of Final Score ${s}_{i}$: Table 3 compares key score metrics for anomaly detection: unimodal scores $s^{m}_{i}$ ($m\in\{P,R\}$) (Euclidean distances, see Section 3.2), scaling factors $w^{m}_{i,0}$ from LSPN in Eq. (2), and our fused score ${s}_{i}$ in Eq. (11). As evidenced in Table 3, unimodal anomaly scores $s^{m}_{i}$ demonstrate suboptimal performance. This limitation stems from the inherent information incompleteness of individual modalities. Our fused score ${s}_{i}$ overcomes these constraints through crossmodal complementarity, achieving significant improvements over unimodal baselines. These quantitative gains are also supported by Figs. 5 and 7.
While $w^{m}_{i,0}$ characterizes the local directions of normal features within local spaces (Section 3.4), its exclusion of the valuable distances, i.e., $s^{m}_{i}$, leads to critical limitations. Specifically, proximity to memory prototypes ($s^{m}_{i}\rightarrow 0$) inherently indicates normality regardless of directional deviations, which is a prior ignored by $w^{m}_{i,0}$. As evidenced in Fig. 7, the score maps obtained by $w^{m}_{i,0}$ exhibit noticeable false positives. Using ${s}_{i}$ in our G2SF mitigates this issue, which benefits from sufficiently leveraging complementary directional and distance information. The comprehensive results are provided in the supplementary material.
@@
Table 5: Investigation on different score aggregation strategies on MVTec 3D-AD dataset.
Ablation of Loss Components: Table 4 systematically evaluates the contribution of each loss component through stepwise removal. Using the separation loss $\mathcal{L}_{sep}$ as the baseline for the metric learning, the addition of the consistency loss $\mathcal{L}_{cns}$ drives an 8.2% I-AUROC improvement, surpassing M3DM [32] across all metrics and achieving parity with CFM [11]. The margin loss $\mathcal{L}_{mar}$ specifically improves pixel-wise localization (AUPRO@30% +0.3%), while the scaling loss $\mathcal{L}_{sc}$ and cross-modal alignment loss $\mathcal{L}_{cma}$ further enhances image-level detection (I-AUROC +1.1/1.0%). These synergistic effects collectively establish G2SF’s state-of-the-art performance, validating the necessity of all proposed loss components.
Investigation of Score Aggregation Strategy: Section 3.6 employs a $\min$ operator in Eq. (11) to aggregate metrics $\{l_{i,j}\}$ from adjacent local spaces. We systematically evaluate three alternatives in Table 5: (1) single-space baseline $l_{i,0}$. (2) $\max$ operator. (3) $\mathrm{mean}$ operator. The $\min$-based strategy achieves superior overall performance (+0.8%/0.7% I-AUROC/AUPRO@1% vs $l_{i,0}$), validating the necessity of multi-space consensus. While $\max$ and $\mathrm{mean}$ operators incorporate adjacent local spaces, they indiscriminately weight less discriminative $l_{i,j}$ from distant regions, compromising detection precision. The $\min$ operator in Eq. (11) prioritizes the most representative local metric through a proper geometric prior as discussed in Section 3.6. The results of all categries of MVTec-3D AD are summarized in the supplementary material.
Method AST [28] M3DM [32] CFM[11] Ours
Frame Rate 4.97 2.12 21.8 6.04
Table 6: Frame rate comparison on MVTec 3D-AD dataset.
Computational Time: Table 6 shows that G2SF detects 6.04 samples per second, sufficient for general industrial requirements.
Analyses of Memory Prototypes and Global Scaling Factors $\sigma_{m}$: Supplementary material provides the sensitivity study about the number of memory prototypes, demonstrating relatively stable performance. Additionally, the investigation on the role of $\sigma_{m}$ is included, showing the necessity of learning optimal $\sigma_{m}$ instead of fixed ones.
5 Conclusion
In this paper, we propose G2SF for multimodal anomaly detection, which learns a unified anisotropic local distance metric to replace the indiscriminative isotropic Euclidean metrics. This metric is formulated by the direction-aware scaling factors predicted by LSPN and original Euclidean distances in individual modalities. We propose new loss functions to enable robust metric learning, serving for discrimination or geometric property maintenance. A geometric prior-induced anomaly scoring strategy is proposed to enhance anomaly detection and segmentation. Extensive experimental results show our state-of-art performance compared to existing competitive methods.
6 Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 72371219 and No. 72001139), and the Guangzhou Municipal Science and Technology Project (No. 2025A04J5288). Corresponding author is Juan Du.
Supplementary Material
@@ B Implementation Details
B.1 Feature Extraction
RGB features are extracted using DINO ViT-B/8 [7] pretrained on ImageNet [13], generating 28×28×76828$\times$ 28$\times$ 76828 × 28 × 768 feature maps upsampled to 56×56×76856$\times$ 56$\times$ 76856 × 56 × 768. Point clouds are processed via Point-MAE [24] pretrained on ShapeNet [8], grouping 3D points into 1024 clusters (feature dimension 1152) and interpolated to 56×56×115256$\times$ 56$\times$ 115256 × 56 × 1152 for spatial alignment with image feature maps. We use PatchCore [27] to construct modality-specific memory banks with 10% coreset sampling. Notably, we consider the original 28×28×76828$\times$ 28$\times$ 76828 × 28 × 768 RGB feature maps instead of the upsampled ones when constructing the memory bank of RGB modality to reduce the memory storage. The above procedure remains the same as done in M3DM [32].
B.2 Data Processing
To accelerate training and underscore discriminative foreground features, we pre-collect features from 800 synthetic anomalous samples per class (Section 3.7) (1000 for Eyecandies), forming a feature dataset equally split for training and validation. The features originate from foreground regions identified via RANSAC plane fitting on 3D point clouds, where areas exceeding 0.005 distance from the fitted plane are retained following [32, 11]. During evaluation on MVTec 3D-AD, background regions remain included. As background features are excluded during training, we set their scaling factors to unity ($w^{m}_{i,0}=1$) in Eq. (3), serving as a linear combination of the original unimodal results. For Eyecandies, background regions are neglected following CFM [11].
Dataset $\alpha$ $\beta$ $\gamma$ $\mu$ Batch Size
MVTec 3D-AD 10 60 8 20 8192
Eyecandies 10 100 2 0.5 15360
Table 1: Turning parameters and batch size in our experiment.
B.3 Parameter Setting
For $\mathcal{L}_{cns}$ in Eq. (6), we set $k=5$ and $\eta_{0}=1.2$. We train our LSPN using an Adam optimizer [21] with learning rate $1.5e^{-4}$ and weight decay $1.5e^{-4}$, combined with $l_{1}$ regularization of weight $1e^{-4}$ and drop out with ratio 0.5 over 80 epochs. Global scaling factors $\sigma^{m}$ are separately optimized at a higher learning rate of $5e^{-3}$ to accelerate convergence. We initialize LSPN by the default strategy in Pytorch, that is, Kaiming Initialization [18]. We observe $w^{m}_{i,j}\approx 1$ under this strategy, meeting the requirement in Section 3.4. For modality $m$, we normalize all Euclidean distances $\{s^{m}_{i,j}\}$ by their mean over the training dataset. Then, we initialize $\sigma^{m}=0.5$ with equal importance from each modality. The tuning parameters about loss components and batch size for MVTec 3D-AD and Eyecandies datasets are listed in Table 1. Finally, we implement our G2SF by Pytorch and run all experiments on Linux with a Nvidia RTX 4090 GPU.
@@ C Learning Curve
Figure 3: Performance over different training epochs on MVTec 3D-AD dataset.
@@
We train multiple models that stop at {2,4,6,8,10,20,…,80} epochs respectively, and report their performances (I-AUROC, P-AUROC, AUPRO@30%, and AUPRO@1%) on MVTec-3D AD in Fig. 3. After only 4 epochs, G2SF outperforms M3DM [32] (94.5% I-AUROC, 99.2%P-AUROC, 96.4% AUPRO@30%, and 39.4%AUPRO@1%). This benefits from our design, as our model evolves from an established Euclidean metric to the target anisotropic metric, see Section 3.4.
Figure 4: Performance over different ratios of coreset sampling on MVTec 3D-AD dataset.
D Sensitivity Study on Sizes of Memory Banks
Since the learned metric $l(\cdot)$ is defined in terms of memory prototypes, this subsection presents a sensitivity study of the anomaly detection performance with respect to the number of prototypes. We vary the coreset sampling ratio (see Section B.1) at {10,7.5,5,2.5,1}levels. At the default 10%, the average numbers of point cloud and RGB image memory prototypes across all categories on the MVTec-3D AD dataset are 83,292 and 20,823, respectively. The corresponding results for the evaluation metrics are summarized in Fig. 4. At the 2.5 % level, G2SF achieves 96.6% I-AUROC, 99.6% P-AUROC, and 97.6% AUPRO@30%. Nevertheless, Fig. 4 demonstrates that the performance of G2SF is relatively stable to the sizes of memory banks.
Metric MVTec-3D AD Eyecandies
I-AUROC P-AUROC AUPRO@30% AUPRO@1% I-AUROC P-AUROC AUPRO@30% AUPRO@1%
Learn 0.971 0.997 0.979 0.468 0.902 0.982 0.898 0.357
$\sigma^{P}$=$\sigma^{R}$=0.5 0.964 0.994 0.972 0.460 0.827 0.977 0.892 0.328
Table 2: Ablation study of global scaling factors on MVTec-3D AD and Eyecandies datasets.
E Analysis on Global Scaling Factors
G2SF defines trainable global scaling factors $\sigma^{m}$ ($m\in{P,R}$) in Eq. (3) to formulate the metric $l(\cdot)$, where $\sigma^{m}$ controls the importance of modality $m$ to final $l(\cdot)$ and normalizes the original Euclidean distances. Fig. 5 demonstrates this mechanism prioritizing image modality over less discriminative point cloud modality of the ChocolateCookie category in Eyecandies dataset, compared with fusion and image score maps. Moreover, ignoring the contribution of the RGB image modality by setting $\sigma^{P}=1$ and $\sigma^{R}=0$ yields significantly degraded results. Furthermore, Table 2 compares performance using learned $\sigma^{m}$ against fixed $\sigma^{P}=\sigma^{R}=0.5$ during inference on both the MVTec 3D-AD and Eyecandies datasets. These results collectively underscore the importance of learning optimal $\sigma^{m}$ during training.
Figure 5: Visualization of learned $\sigma^{m}$ and fixed $\sigma^{P}=1$ for an example from ChocolateCookie category in Eyecandies dataset.
F Results on Eyecandies
We provide the quantitative results for all categories in Eyecandies in Table 4. G2SF achieves the best performance of all metrics in terms of +0.5% I-AUROC and +0.8% AUPRO@30% over 2M3DF [1], +0.5% P-AUROC over M3DM [32], +2.2%AUPRO@1%, +2.4%AUPRO@10%, and +2.8%AUPRO@5% over CFM [11]. The qualitative results are illustrated in Fig. 6.
Figure 6: Qualitative results on Eyecandies dataset.
Method Can. C. Cho. C. Cho. P. Conf. Gum. B. Haz. T. Lic. S. Lollip. Marsh. Pep. C. Mean
I-AUROC AST[28] 0.574 0.747 0.747 0.889 0.596 0.617 0.816 0.841 0.987 0.987 0.780
M3DM* [32] 0.597 0.954 0.931 0.990 0.883 0.666 0.923 0.888 0.995 1.000 0.882
CFM [11] 0.680 0.931 0.952 0.880 0.865 0.782 0.917 0.840 0.998 0.962 0.881
2M3DF [1] 0.753 0.955 0.937 0.967 0.901 0.792 0.889 0.913 0.980 0.893 0.897
Ours 0.696 0.963 0.967 1.000 0.907 0.701 0.939 0.855 0.989 1.000 0.902
P-AUROC AST[28] 0.763 0.960 0.911 0.969 0.788 0.837 0.918 0.924 0.983 0.968 0.902
M3DM* [32] 0.968 0.986 0.964 0.998 0.976 0.928 0.976 0.988 0.996 0.995 0.977
CFM [11] 0.985 0.984 0.961 0.986 0.958 0.937 0.968 0.981 0.994 0.978 0.973
LSFA [30] 0.969 0.957 0.967 0.996 0.971 0.938 0.970 0.990 0.998 0.987 0.974
Ours 0.981 0.982 0.977 0.998 0.982 0.936 0.981 0.989 0.997 0.998 0.982
AUPRO@30% AST[28] 0.514 0.835 0.714 0.905 0.587 0.590 0.736 0.769 0.918 0.878 0.744
M3DM* [32] 0.889 0.921 0.808 0.982 0.889 0.675 0.872 0.901 0.964 0.973 0.887
CFM [11] 0.942 0.902 0.831 0.965 0.875 0.762 0.791 0.913 0.939 0.949 0.887
2M3DF [1] 0.924 0.935 0.820 0.940 0.875 0.781 0.816 0.923 0.958 0.926 0.890
Ours 0.928 0.897 0.843 0.982 0.890 0.687 0.888 0.913 0.971 0.981 0.898
AUPRO@1% AST[28] 0.035 0.230 0.129 0.234 0.092 0.069 0.139 0.090 0.255 0.224 0.149
M3DM* [32] 0.166 0.388 0.329 0.486 0.315 0.131 0.323 0.258 0.462 0.454 0.331
CFM [11] 0.229 0.397 0.345 0.389 0.353 0.188 0.333 0.236 0.455 0.428 0.335
Ours 0.174 0.416 0.377 0.487 0.360 0.164 0.353 0.281 0.478 0.479 0.357
AUPRO@10% AST[28] 0.285 0.709 0.545 0.770 0.404 0.350 0.584 0.544 0.770 0.744 0.570
M3DM* [32] 0.677 0.836 0.698 0.947 0.754 0.410 0.732 0.712 0.913 0.924 0.760
CFM [11] 0.827 0.815 0.731 0.896 0.741 0.550 0.663 0.739 0.893 0.868 0.772
Ours 0.784 0.820 0.752 0.946 0.780 0.490 0.787 0.739 0.922 0.943 0.796
AUPRO@5% AST[28] 0.173 0.592 0.421 0.635 0.288 0.242 0.461 0.378 0.634 0.617 0.444
M3DM* [32] 0.479 0.759 0.626 0.894 0.655 0.300 0.634 0.562 0.849 0.861 0.661
CFM [11] 0.662 0.750 0.653 0.801 0.657 0.427 0.609 0.552 0.838 0.796 0.675
Ours 0.578 0.759 0.687 0.892 0.703 0.399 0.698 0.568 0.858 0.889 0.703
Table 3: Results on Eyecandies. Best results in bold, runner-ups underlined.
G Additional Results on MVTec-3D AD
@@
We summarize the results of different score metrics for all categories in MVTec-3D AD in Table 5. The results demonstrate that the final score $s_{i}$ outperforms unimodal scores $\{s^{m}_{i,0}\}$ and scaling factors $\{w^{m}_{i,0}\}$ almost all categories. Additionally, comprehensive quantitative results of various score aggregation operators are provided in 6, which validate the superior performance of $\min$ operator over the majority of classes, owing to its geometric interpretation, as discussed in Section 3.6.
Furthermore, Fig. 7 provides additional qualitative results for all classes in MVTec-3D AD. Consistent with Fig. 5, our G2SF suppresses anomaly scores of normal regions and provides clearer anomaly boundaries, compared to M3DM [32] and CFM [11].
AUPRO@1% BTF[19] 0.428 0.365 0.452 0.431 0.370 0.244 0.427 0.470 0.298 0.345 0.383
AST[28] 0.388 0.322 0.470 0.411 0.328 0.275 0.474 0.487 0.360 0.474 0.398
Shape-guided[10] - - - - - - - - - - 0.456
M3DM [32] 0.414 0.395 0.447 0.318 0.422 0.335 0.444 0.351 0.416 0.398 0.394
CFM [11] 0.459 0.431 0.485 0.469 0.394 0.413 0.468 0.487 0.464 0.476 0.455
Ours 0.481 0.443 0.484 0.471 0.410 0.468 0.487 0.499 0.468 0.471 0.468
.@10% CFM [11] 0.937 0.917 0.947 0.897 0.855 0.906 0.942 0.947 0.926 0.944 0.922
Ours 0.946 0.928 0.947 0.937 0.912 0.932 0.946 0.950 0.933 0.942 0.937
.@5% CFM [11] 0.877 0.843 0.894 0.840 0.765 0.828 0.884 0.894 0.865 0.889 0.858
Ours 0.892 0.860 0.893 0.877 0.828 0.873 0.892 0.899 0.870 0.885 0.877
@@
Table 4: Additional AUPRO@1%, AUPRO@10% and AUPRO@5% results on MVTec-3D AD. Best results in bold, runner-ups underlined.
@@
Table 5: Results of various score metrics on MVTec-3D AD. Best results in bold, runner-ups underlined.
@@
Table 6: Results of various score aggregation strategies on MVTec-3D AD. Best results in bold, runner-ups underlined.
Figure 7: Additional qualitative results on MVTec-3D AD dataset.